OpenStudy (anonymous):

what is the volume of a triangular prism that has a base area of 30 mm2 and a height of 18 mm? a rectangle has a length of 6 meters and a perimeter of 20 meters. What is the width of the rectangle?

hartnn (hartnn):

For first question the formula is VOLUME=BASE AREA * HEIGHT

OpenStudy (anonymous):

For the second one, I'll give you a hint. Perimeter=2(length)+2(width) In this case, 20=2(6)+2w

OpenStudy (anonymous):

find w.

OpenStudy (anonymous):

Triangular prism: 30 * 18 = 540 mm^3

OpenStudy (anonymous):

the volume is 270mm3 -volume of a triangular prism=1/2xlengthxwidthxheight -area of the triangular base=1/2xbxh the width is 4 -to find perimeter you add all the sides -if 6 is length, their are two sides that have to be 6 -that=12 -20-12=8 -8/2=4

OpenStudy (anonymous):

Perimeter: 20 - 12 = 8 The width is 4

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

np (: